As you work through the parts of this question you are going to show that the Maxwell equations naturally contain electromagnetic waves. In a region of space that is void of all charges and currents, ρ = 0 and J~ = 0 the Maxwell equations come out to be:Question 1: As you work through the parts of this question you are going to show that the Maxwell equations naturally contain
Answer #1

a) In the region of space that lacks charges and currents, \rho =0 and J=0, the Maxwell's equations can be written as,

\bigtriangledown.E=0 (Gauss law in electrostatics) ---------(1)

\bigtriangledown.B=0 (Gauss law in magnetostatics) -----------(2)

\bigtriangledown\timesE=-\frac{\partial B}{\partial t} (Faradays law) ---------(3)

\bigtriangledown\timesB=\muo\epsilono\frac{\partial E}{\partial t} (Ampere-Maxwell law) ------------(4)

a) Taking curl of both sides of equation (4),

\bigtriangledown\times(\bigtriangledown\timesB)=\muo\epsilono\bigtriangledown\times\frac{\partial E}{\partial t}

\bigtriangledown(\bigtriangledown.B)-\bigtriangledown2B=-\muo\epsilono\frac{\partial^2 B}{\partial t^2} (from 3)

Using (2),\bigtriangledown2B=\muo\epsilono\frac{\partial^2 B}{\partial t^2} --------(5)

where c=(1/\muo\epsilono)1/2

This is the wave equation for magnetic field.

b) The wave equation can be written as a function that is travelling in the direction of the wave and the other travelling in the opposite direction.

Let B=B1 cos(kx-\omegat)+B2 cos(kx+\omegat)

where B1 and B2 are constant vectors

\bigtriangledown2B=-k2 (B1 cos(kx-\omegat)+B2 cos(kx+\omega​​​​​​​t))

\muo\epsilono\frac{\partial^2 B}{\partial t^2}=-\muo\epsilono\omega2(B1 cos(kx-\omegat)+B2 cos(kx+\omega​​​​​​​t)

Subsituting in (5),

k2=\muo\epsilono\omega2 which is true, and hence it satisfies the wave equation

Speed of the wave c=(1/\muo\epsilono)1/2

Thus, from the above equation c=\omega/k which is the speed of these waves

c) Given, f=f1(kx-\omegat)+f2(kx+\omega​​​​​​​t), where f1 and f2 are generic vector field functions

Substituting this into wave equation similar to equation (5)

\bigtriangledown2f=\muo\epsilono0 f Ꭷt2

We get k2=\muo\epsilono\omega2 , which is a solution of wave equation.

thus, any function that comes with these kinds of variable dependence are solutions to the wave equations and are thus waves. Thus, for plane waves f1 and f2 are costants. For spherical waves, they depend inversely on r. Thus, the amplitudes of f1 and f2 give the shape of the wave.